# Unexpected properties of Markov chains with an infinitely countable set of states

### v4

@@ 4. Counterexamples
$\left(\begin{matrix}0&1&0&0&0&0&0&0&\ldots\\ \epsilon&0&1-\epsilon&0&0&0&0&0&\ldots\\ \frac{1}{3}-\epsilon&\epsilon&0&\epsilon&\frac{2}{3}-\epsilon&0&0&0&\ldots\\ 0&\frac{1}{2}+\epsilon&0&0&0&\frac{1}{2}-\epsilon&0&0&\ldots\\ 0&0&\frac{1}{3}-\epsilon&\epsilon&0&\epsilon&\frac{2}{3}-\epsilon&0&\ldots\\ 0&0&0&\frac{1}{2}+\epsilon&0&0&0&\frac{1}{2}-\epsilon&\ldots\\ \vdots&\vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\ddots&\ddots\end{matrix}\right),$
@@
$\left(\begin{matrix}0&1&0&0&0&0&0&0&\ldots\\ 1-\epsilon&0&\epsilon&0&0&0&0&0&\ldots\\ \frac{2}{3}-\epsilon&\epsilon&0&\epsilon&\frac{1}{3}-\epsilon&0&0&0&\ldots\\ 0&\frac{1}{2}-\epsilon&0&0&0&\frac{1}{2}+\epsilon&0&0&\ldots\\ 0&0&\frac{2}{3}-\epsilon&\epsilon&0&\epsilon&\frac{1}{3}-\epsilon&0&\ldots\\ 0&0&0&\frac{1}{2}-\epsilon&0&0&0&\frac{1}{2}+\epsilon&\ldots\\ \vdots&\vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\ddots&\ddots\end{matrix}\right),$
@@ 5. Discussion
No conflict of interest was reported by the author

### v5

@@ 4. Counterexamples
$\left(\begin{matrix}0&1&0&0&0&0&0&0&0&\ldots\\ \epsilon&0&1-\epsilon&0&0&0&0&0&0&\ldots\\ \frac{1}{3}-\epsilon&\epsilon&0&\epsilon&\frac{2}{3}-\epsilon&0&0&0&0&\ldots\\ 0&\frac{1}{2}+\epsilon&0&0&0&\frac{1}{2}-\epsilon&0&0&0&\ldots\\ 0&0&\frac{1}{3}-\epsilon&\epsilon&0&\epsilon&\frac{2}{3}-\epsilon&0&0&\ldots\\ 0&0&0&\frac{1}{2}+\epsilon&0&0&0&\frac{1}{2}-\epsilon&0&\ldots\\ \vdots&\vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\end{matrix}\right),$
@@
$\left(\begin{matrix}0&1&0&0&0&0&0&0&0&\ldots\\ 1-\epsilon&0&\epsilon&0&0&0&0&0&0&\ldots\\ \frac{2}{3}-\epsilon&\epsilon&0&\epsilon&\frac{1}{3}-\epsilon&0&0&0&0&\ldots\\ 0&\frac{1}{2}-\epsilon&0&0&0&\frac{1}{2}+\epsilon&0&0&0&\ldots\\ 0&0&\frac{2}{3}-\epsilon&\epsilon&0&\epsilon&\frac{1}{3}-\epsilon&0&0&\ldots\\ 0&0&0&\frac{1}{2}-\epsilon&0&0&0&\frac{1}{2}+\epsilon&0&\ldots\\ \vdots&\vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\end{matrix}\right),$
@@ 5. Discussion
In our opinion, the following future study might be interesting.
1. The main result of this paper assumes that the matrix of transition probabilities form connected domain. However, for many transition probability matrices that do not form connected domain the results remain true, and the proof provided in this paper does not use this class of matrices explicitly. Can the class of transition probability matrices be extended?
2. One of the conditions of Theorem 1.2 requires that the diagonal elements of the matrix satisfy the condition: $1-p_{i,i}>\epsilon$ for all $i\geq 0$, where $\epsilon$ is some positive value. Can this condition be relaxed in some way? Say, one can assume that $p_{i,i}$ tends to $1$ as $i\to\infty$, but at some slow rate that will enable us to claim that the main result still holds true.
@@
No conflict of interest was reported by the author.